OpenStudy (anonymous):

An airplane flies 200km due west from city A to city B and then 300 km in the direction of 30.0 degrees north of west from city B to city C.?

OpenStudy (anonymous):

If you need to find the length from city A to C, then you can make a triangle. Here you have the length from A to B = 200, and from B to C = 300. And the angle B would be: \[180-30=150\] Now you can find the length from A to C using the law of cosines: \[c^2=a^2+b^2-2*a*b*\cos(B)\]